OpenStudy (annalee607):

Two square pyramids have the same volume. For the first pyramid , the side length of the base is 14 inches, and the height is 24 inches. The second pyramid has a height of 96 inches. What is the side length of the base of the second pyramid.

Vocaloid (vocaloid):

set up 2 volume equations

Vocaloid (vocaloid):

Volume = side length^2 * height * (1/3) so (14)(14)(24)(1/3) = (96)(1/3)x^2

Vocaloid (vocaloid):

then the 1/3's cancel out (14)(14)(24) = 96x^2 solve for x

OpenStudy (annalee607):

4704=9216x

Vocaloid (vocaloid):

nope not quite

Vocaloid (vocaloid):

the ^2 only applies to the x so 4704 = 96x^2 now solve for x

OpenStudy (annalee607):

48?

Vocaloid (vocaloid):

nope what's 4704/96?

OpenStudy (annalee607):

49

Vocaloid (vocaloid):

good. now what's the square root of 49?

OpenStudy (annalee607):

7

Vocaloid (vocaloid):

good, that's your final answer